OpenStudy (star):

f(x) = (a/x) + ((bsinx)/(x)) + ((cx)/(pi)) + ((dx^2)/(pi^2)) what is lim x-> 0 for f(x)? (i'll type the equation in the formula editor now to make it look clearer)

OpenStudy (osanseviero):

yes please

OpenStudy (anonymous):

use \frac{top}{bottom} for fractions

OpenStudy (star):

\[\lim_{x \rightarrow 0} \frac{a}{x} + \frac{bsinx}{x} + \frac{cx}{\pi} + \frac{dx^2}{\pi^2}\]

OpenStudy (star):

thank you for the code i needed it! was looking for it everywhere

OpenStudy (zarkon):

what is a?

OpenStudy (anonymous):

probably a positive integer

OpenStudy (star):

where a, b, c, d is an unknown value

OpenStudy (star):

the answer can be in terms of a, b, c, d

OpenStudy (zarkon):

the limit DNE if \[a\neq0\]

OpenStudy (zarkon):

if a=0 the answer is b

OpenStudy (zarkon):

I'm just taking the limit

myininaya (myininaya):

yes i agree a has to be 0 for the limit to exist and if a=0, then the limit is b i agree zarkon gj :)

OpenStudy (anonymous):

oh i though you said L=0

OpenStudy (star):

may i ask why a = 0? sorry :(

OpenStudy (anonymous):

cannot equal 0

OpenStudy (anonymous):

can b = 0?

OpenStudy (star):

oops i meant \[\neq\]

OpenStudy (star):

i'm really confused.. so in order for the limit to exist a cannot be 0 and therefore b is 0?

myininaya (myininaya):

\[\lim(b\cdot \frac{\sin(x)}{x})+\lim(\frac{a \pi^2+cx^2\pi+dx^3}{\pi^2x})\] \[=b(1)+\lim \frac{a \pi^2+cx^2\pi+dx^3}{\pi^2x}\] so for that second limit to exist we need it to be 0/0 when we plug in 0 but we get \[\frac{a \pi^2}{0}\] s0 this implies a=0 in order for limit to exist and if a=0 we have \[=b+\lim \frac{cx^2 \pi+dx^3}{\pi^2 x}=b+\lim \frac{2xc \pi +3dx^2}{\pi^2}=b+\frac{0}{\pi^2}=b\]

OpenStudy (star):

thank you so much! the answer's perfect :) thank youuu!!